OpenStudy (anonymous):

How do I write the equation for the line through (4,7) and -2,1) in slope-intercept form?

OpenStudy (dan815):

hey

OpenStudy (dan815):

both of these 2 points belong on the same line, a line has constant slope , which means the same ratio of rise over run

OpenStudy (dan815):

y=mx+b \[given~~points~~p1~and~p2~where~p1=(x1,y1)~~and~p2=(x2,y2)\\ y-y1=(x-x1) \\m=\frac{y1-y2}{x1-x2}\\ you~can~rearrange~into~y=mx+b, where~b~is~a~constant\\~and~its~also~known~as~ the~y-intercept\] as the y-intercept happens when x=0 so y=m(0)+b, y=b when x=0 so b is the yintercept

OpenStudy (dan815):

oh the equation of the line in there is y-y1=m(x-x1)

OpenStudy (dan815):

we rearrange that into slope intercept form

hero (hero):

@Poohbear_Sophmore do you know how to find the slope of the line given two points?

OpenStudy (anonymous):

@Hero Um I guess not? Should I graph it first to find the slope?